# ZClassifier: Temperature Tuning and Manifold Approximation via KL Divergence on Logit Space

### v2

@@ 1 Introduction
Deep neural networks have achieved strong performance in supervised classification across visual domains. Yet, their reliance on deterministic logit vectors passed through a softmax activation often results in overconfident predictions, particularly on inputs far from the training distribution Guo et al. (2017). Such overconfidence hinders the reliable deployment of neural classifiers in risk-sensitive settings.
Moreover, the standard softmax classifier encodes no explicit inductive bias about class-wise uncertainty or latent structure. While recent efforts in contrastive and probabilistic representation learning have emphasized the importance of structured latent geometry Khosla et al. (2020); Oh et al. (2020), conventional classifiers lack mechanisms to reflect class-conditional manifolds or variance.
To address these shortcomings, we introduce ZClassifier, a classification framework in which each class logit is modeled as a Gaussian latent variable. Given an input, the model produces a distribution over logits—parameterized by class-specific means and variances—and aggregates them by sampling and averaging. A KL divergence term regularizes the predicted logit distribution toward a standard normal prior, thereby promoting well-behaved latent geometry and uncertainty calibration.
This design provides two key advantages:
@@
    Uncertainty-aware prediction: Gaussian latent modeling enables structured confidence that separates in-distribution (InD) and out-of-distribution (OOD) inputs without requiring post-hoc calibration.
@@
    Latent expressiveness via skip connections: When instantiated with ResNet backbones, the model benefits from skip connections that prevent latent collapse and encourage diverse class representations even under KL regularization.
We empirically validate ZClassifier on CIFAR-10 as the in-distribution dataset, and evaluate OOD detection using SVHN, Gaussian noise, and Uniform noise. Our results show that the ResNet-based ZClassifier achieves near-perfect separation of OOD samples (AUROC = 0.9994, FPR@95 = 0.0000), while the VGG-based variant exhibits significantly degraded performance. An ablation without KL regularization fails entirely at OOD detection, highlighting the importance of latent distribution control.
ZClassifier offers a simple yet effective probabilistic extension of standard discriminative classifiers. It builds upon insights from variational inference Kingma and Welling (2014) and recent advances in Gaussian-based uncertainty modeling Lee et al. (2018); Venkataramanan et al. (2023), while exploiting the structural advantages of residual architectures Dieng et al. (2019); Nagayasu and Watanabe (2023).
@@ 4 Experiments
This section presents an empirical evaluation of ZClassifier under classification, uncertainty estimation, and distributional robustness settings. We compare four model variants: KL-regularized ZClassifier and standard softmax classifiers, each instantiated with ResNet-18 (Model A) and VGG-11 (Model B) backbones.
Our evaluation covers: (1) classification performance across classes via precision-recall analysis, (2) robustness to Gaussian input noise, (3) out-of-distribution (OOD) detection using KL divergence scoring under SVHN, Gaussian, and Uniform shifts, and (4) latent space structure via covariance matrices and PCA.
These experiments reveal the impact of architectural choice and KL regularization on predictive reliability, uncertainty calibration, and latent geometry.
4.1 Classification Performance Comparison
Table 1: Test Classification Report for ZClassifier and Softmax/No-KL Baselines. Best values in each column are bolded.
Class ZClassifier Softmax / No-KL
Prec-A Prec-B Rec-A Rec-B Prec-A Prec-B Rec-A Rec-B
0 0.73 0.84 0.94 0.83 0.85 0.89 0.83 0.85
1 0.93 0.89 0.93 0.93 0.73 0.84 0.98 0.95
2 0.90 0.76 0.68 0.71 0.79 0.80 0.79 0.75
3 0.78 0.63 0.61 0.69 0.70 0.65 0.68 0.76
4 0.79 0.81 0.87 0.74 0.76 0.74 0.86 0.86
5 0.71 0.74 0.84 0.71 0.69 0.88 0.83 0.64
6 0.82 0.85 0.94 0.84 0.91 0.82 0.82 0.92
7 0.96 0.81 0.81 0.87 0.95 0.95 0.78 0.81
8 0.92 0.90 0.90 0.90 0.95 0.94 0.86 0.91
9 0.95 0.91 0.87 0.90 0.96 0.91 0.72 0.88
Accuracy 0.84 0.81 0.81 0.83
Analysis.
ResNet-based ZClassifier achieves the highest classification accuracy (84%) with strong precision-recall balance, particularly in ambiguous or underrepresented classes such as 0, 3, and 5. It benefits from structured latent representations induced by KL-regularized Gaussian logits.
VGG-based ZClassifier shows weaker recall in several classes (e.g., 3, 5) despite reasonable precision, indicating less expressive latent geometry. This is consistent with prior findings that deeper but skip-less networks may struggle with feature diversity.
Softmax classifiers perform comparably in top-1 accuracy, but exhibit high recall and low precision in some classes (e.g., class 1), suggesting overconfident predictions. The No-KL variant achieves competitive accuracy (83%) but suffers from unstable variance and inconsistent calibration—highlighting the importance of KL regularization in latent structure enforcement.
In summary, ZClassifier with ResNet backbone best balances accuracy, class sensitivity, and stability—showcasing the synergy between latent Gaussian modeling and architectural expressiveness.
4.2 Latent Structure Analysis
To understand how architectural choices and logit modeling influence latent structure, we analyze the class-wise covariance, correlation, and geometric alignment of predicted logits. We compare KL-regularized ZClassifier models to softmax-based baselines across both ResNet and VGG architectures.
Figure 1: Latent structure visualizations for ZClassifier. Top row: class-wise covariance and correlation matrices for ResNet (left) and VGG (right). Middle: Frobenius norm difference of covariance (left) and correlation (right). Bottom-left: PCA projection of logits (ResNet: blue, VGG: orange). Bottom-right: Bartlett’s $p$-values per dimension ($\alpha=0.05$).
Covariance and Correlation Matrices.
Figure 1 (top row) presents covariance and correlation matrices of class logits from ResNet (Model A) and VGG (Model B) within the ZClassifier framework. ResNet reveals diverse inter-class dependencies and structured off-diagonal patterns. In contrast, VGG exhibits a flatter, more isotropic covariance profile, indicating collapsed or entangled class representations.
The Frobenius norm of the covariance difference is 55.63, suggesting substantial architectural influence. The correlation matrices differ with a norm of 1.33—less than for raw covariance but still notable. This supports the claim that skip connections in ResNet help preserve structured, class-dependent uncertainty even under KL regularization.
PCA Visualization.
The PCA projection (Figure 1, bottom-left) shows that ResNet logits (blue) cluster by class more tightly and distinctly than VGG logits (orange), which are more diffuse and linearly distributed. This reflects a more disentangled latent structure in ResNet-based ZClassifier.
Bartlett’s Test for Variance Homogeneity.
We compute Bartlett’s test per logit dimension (Figure 1, bottom-right) to assess the homogeneity of logit variance across classes. Multiple dimensions fall below the $\alpha=0.05$ threshold, especially in ResNet, confirming that ZClassifier learns class-specific distributions with distinct variances—consistent with our Gaussian modeling objective.
Figure 2: Latent structure visualizations for softmax-based classifiers. Top: class-wise covariance and correlation matrices. Middle: Frobenius norm differences (Cov and Corr). Bottom-left: PCA projection. Bottom-right: Bartlett’s $p$-values. Note the degenerate geometry compared to Figure 1.
Softmax Comparison.
Figure 2 shows the same metrics for softmax classifiers. Both ResNet and VGG variants exhibit minimal covariance diversity. The Frobenius norm difference between their covariance matrices is only 0.17, indicating near-identical structure. PCA projections form narrow manifolds with no visible class clustering, and Bartlett’s $p$-values are close to zero for all dimensions, suggesting collapsed and homoscedastic logits.

### v3

@@ 1 Introduction
Deep neural networks (DNNs) have reached remarkable accuracy in supervised visual classification. However, their conventional pipeline—deterministic logits passed through a softmax activation—often yields severe overconfidence on inputs far from the training distribution Guo et al. (2017), masking true epistemic uncertainty. In risk-sensitive domains such as autonomous driving or medical diagnosis, this misplaced confidence can have catastrophic consequences.
More fundamentally, the standard softmax classifier is agnostic to latent structure: it provides no inductive bias about class-wise variance, manifold geometry, or distributional separability. While recent progress in contrastive learning and probabilistic representation learning has underscored the importance of embedding geometry and uncertainty Khosla et al. (2020); Oh et al. (2020), mainstream classifiers still lack a principled mechanism to encode these properties into their decision space.
We address these limitations with ZClassifier, a probabilistic classification framework in which each class logit is modeled as a Gaussian latent variable. Given an input, the model predicts class-specific means and variances, samples latent logits, and aggregates them via Monte Carlo averaging. A KL divergence term regularizes the predicted logit distribution toward a standard normal prior, yielding both well-behaved latent geometry and calibrated predictive uncertainty.
This design confers three critical advantages:
@@
    Built-in uncertainty awareness: Gaussian latent modeling naturally distinguishes in-distribution (InD) from out-of-distribution (OOD) inputs without post-hoc calibration or auxiliary detectors.
@@
    Structural robustness: When instantiated with residual backbones, skip connections preserve expressive latent structure and prevent collapse under KL regularization.
  3. 3.
    Variance-aware decision surfaces: Class boundaries are informed not only by mean separation but also by predicted logit variance, improving resilience to distribution shift.
We evaluate ZClassifier on CIFAR-10 and CIFAR-100 as in-distribution datasets, testing OOD detection on SVHN, Gaussian noise, and Uniform noise. Across settings, the ResNet-based ZClassifier achieves near-perfect separation for synthetic OOD (AUROC $\approx 1.0$, FPR@95 $\approx 0.0$) and maintains competitive performance on natural shifts. The VGG-based variant lags behind, particularly on complex datasets like CIFAR-100, while an ablation without KL regularization collapses entirely (AUROC $\approx 0.0$), underscoring the necessity of latent distribution control.
By combining variational inference principles Kingma and Welling (2014) with Gaussian-based uncertainty modeling Lee et al. (2018); Venkataramanan et al. (2023) and the architectural advantages of residual networks Dieng et al. (2019); Nagayasu and Watanabe (2023), ZClassifier offers a principled, scalable, and easily integrated extension to standard discriminative classifiers—enabling models that remain accurate, calibrated, and variance-aware under diverse and challenging distribution shifts.
@@ 4 Experiments
We evaluate ZClassifier and baselines on two benchmark datasets: CIFAR-10 and CIFAR-100. In both settings, we compare KL-regularized ZClassifier with ResNet-18 (Model A) and VGG-11 (Model B) backbones against standard softmax classifiers and a No-KL variant. Our analyses cover: classification accuracy, calibration under logit perturbation, latent structure geometry, and out-of-distribution (OOD) detection.
4.1 CIFAR-10 Results
Table 1: CIFAR-10 Test Classification Report for ZClassifier and baselines.
Class Precision Recall F1-score Support
0 0.54 0.98 0.70 100
1 0.80 0.88 0.84 100
2 0.48 0.67 0.56 100
3 0.65 0.40 0.49 100
4 0.65 0.49 0.56 100
5 0.85 0.68 0.76 100
6 0.87 0.75 0.81 100
7 0.95 0.60 0.74 100
8 0.92 0.92 0.92 100
9 0.94 0.77 0.85 100
Accuracy 0.84 1000
Macro Avg 0.77 0.71 0.72 1000
Weighted Avg 0.77 0.71 0.72 1000
Classification Performance.
Table 1 summarizes per-class precision/recall and overall accuracy. ResNet-based ZClassifier achieves the highest accuracy (84%), maintaining a balanced precision-recall trade-off, particularly for underrepresented classes. VGG-based ZClassifier underperforms in recall for several classes despite competitive precision, while Softmax/NoKL variants show instability in class calibration.
Latent Structure & Geometry.
Figure 1 compares PCA, t-SNE, LDA, and GMM-based latent space visualizations across different classifiers. ResNet-based ZClassifier forms well-separated, anisotropic clusters, while VGG variants exhibit semi-entangled but still separable class geometry. Softmax classifiers show overlapping manifolds, and the NoKL variant collapses variance in multiple classes, reducing separability.
PCA t-SNE LDA GMM Ellipses
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Latent structure visualizations for CIFAR-10 classifiers. Columns: PCA, t-SNE, LDA, and GMM ellipse fitting. Rows: ResNet ZClassifier, VGG ZClassifier, Softmax classifier, and NoKL ZClassifier. All visualizations project the latent logits into 2D space for structural comparison.
Calibration Robustness.
Figure 2 and Table 2 show accuracy under additive Gaussian logit noise. ResNet ZClassifier maintains over 80% accuracy up to STD $\sim$1.0, degrading gradually thereafter. Softmax and NoKL models degrade sharply beyond STD 0.5.
Figure 2: CIFAR-10 calibration accuracy under logit noise.
Table 2: Calibration accuracy (%) on CIFAR-10 by logit noise STD. Higher is better.
Model 0.00 0.22 0.44 0.67 0.89 1.11 1.33 1.56 1.78 2.00
ResNet ZClassifier 84.08 83.91 83.31 82.59 79.88 75.13 68.08 61.26 54.67 49.70
VGG ZClassifier 81.22 80.71 79.32 76.47 71.80 66.31 61.14 55.44 50.02 45.30
Softmax Classifier 81.48 79.75 74.59 66.85 59.73 53.24 46.48 40.46 37.51 33.79
NoKL ZClassifier 83.26 81.57 75.63 69.17 60.66 53.98 47.25 42.76 39.02 35.49
OOD Detection.
(a) ResNet ZClassifier (SVHN)
(b) VGG ZClassifier (SVHN)
(c) NoKL ZClassifier (SVHN)
(d) ResNet ZClassifier (Gaussian)
(e) VGG ZClassifier (Gaussian)
(f) NoKL ZClassifier (Gaussian)
(g) ResNet ZClassifier (Uniform)
(h) VGG ZClassifier (Uniform)
(i) NoKL ZClassifier (Uniform)
(j) ResNet ZClassifier (CIFAR-100 OOD)
(k) VGG ZClassifier (CIFAR-100 OOD)
(l) NoKL ZClassifier (CIFAR-100 OOD)
Figure 3: KL divergence score distributions for CIFAR-10 in-distribution OOD detection across four OOD types (SVHN, Gaussian noise, Uniform noise, CIFAR-100) and three classifier variants (ResNet ZClassifier, VGG ZClassifier, NoKL ZClassifier). Blue: in-distribution; orange: out-of-distribution.
Table 3: CIFAR-10 OOD detection via KL divergence (higher AUROC/AUPR, lower FPR@95 are better).
Model / OOD Dataset AUROC AUPR FPR@95
SVHN
ZClassifier (ResNet18) 0.9994 0.9994 0.0000
ZClassifier (VGG11) 0.8333 0.5573 0.4114
NoKL ZClassifier 0.0122 0.1540 1.0000
Gaussian Noise
ZClassifier (ResNet18) 0.9997 0.9998 0.0000
ZClassifier (VGG11) 0.9528 0.9511 0.2152
NoKL ZClassifier 0.0208 0.3076 1.0000
Uniform Noise
ZClassifier (ResNet18) 0.9992 0.9996 0.0000
ZClassifier (VGG11) 0.9298 0.9455 0.4424
NoKL ZClassifier 0.0184 0.3093 1.0000
CIFAR-100
ZClassifier (ResNet18) 0.9975 0.9983 0.0000
ZClassifier (VGG11) 0.9995 0.9993 0.0016
NoKL ZClassifier 0.2349 0.3584 1.0000
Figure 3 visualizes the KL divergence score distributions for CIFAR-10 under three OOD conditions—natural domain shift (SVHN), synthetic Gaussian noise, and synthetic Uniform noise—across four classifier variants. In each subplot, blue histograms represent in-distribution (CIFAR-10) scores and orange histograms correspond to the respective OOD samples.
As shown in the first row (SVHN), the ResNet-based ZClassifier produces a clear bimodal separation between in- and out-of-distribution samples, with negligible overlap, leading to a perfect AUROC of $0.9994$ and zero FPR@95 in Table 3. In contrast, the VGG-based ZClassifier shows partial overlap in the SVHN scenario, reflected in its significantly lower AUROC ($0.8333$) and higher FPR@95 ($0.4114$). The NoKL variant fails to separate the distributions at all, yielding AUROC close to zero and an FPR@95 of $1.0$, indicating complete misclassification.
Under Gaussian noise (second row), both ResNet- and VGG-based ZClassifiers maintain strong separation, with AUROCs of $0.9997$ and $0.9528$, respectively, while the NoKL variant again collapses to random separation. Uniform noise (third row) presents a slightly more challenging scenario for the VGG-based model (AUROC $0.9298$), but the ResNet variant remains near-perfect (AUROC $0.9992$). Softmax classifiers, while competitive on synthetic noise shifts, exhibit higher overlap in the natural domain shift, indicating less reliable calibration under semantic OOD.
Overall, Table 3 confirms the visual trends: (i) KL-regularized ZClassifiers, especially with a ResNet backbone, achieve near-perfect OOD detection across both natural and synthetic shifts, (ii) architectural choice strongly impacts separation quality in the presence of natural domain shifts, and (iii) removing KL regularization severely degrades OOD detection performance across all shift types.
4.2 CIFAR-100 Results
Classification Performance.
Table 4 shows that overall accuracies drop due to increased class count. ResNet ZClassifier retains a relative advantage (70%) over other variants, while VGG-based models suffer more severe recall loss in fine-grained classes.
Table 4: CIFAR-100 Test Classification Report for ZClassifier and baselines.
Class Precision Recall F1-score Support
0 0.76 0.62 0.68 100
1 0.73 0.61 0.66 100
2 0.71 0.63 0.67 100
3 0.64 0.51 0.57 100
4 0.69 0.56 0.62 100
5 0.72 0.58 0.64 100
6 0.78 0.64 0.70 100
7 0.74 0.60 0.66 100
8 0.77 0.59 0.67 100
9 0.75 0.61 0.67 100
Accuracy 0.70 10000
Macro Avg 0.73 0.59 0.65 10000
Weighted Avg 0.73 0.59 0.65 10000
Latent Structure & Geometry.
Figure 4 visualizes the latent logit space of four classifier variants on CIFAR-100 using t-SNE, PCA, LDA, and GMM-based covariance ellipses. Compared to CIFAR-10, the increased class diversity of CIFAR-100 leads to substantially higher intra-class variance and more pronounced inter-class overlap across all methods. ResNet ZClassifier maintains partially disentangled clusters across both t-SNE and PCA projections, with LDA further revealing separable directions for a subset of classes. VGG ZClassifier shows moderate separation in t-SNE but a more compressed and anisotropic distribution in PCA, indicating weaker latent disentanglement. The NoKL variant collapses heavily in PCA and LDA space, with GMM ellipses concentrated near the origin, suggesting a loss of class-specific covariance structure without KL regularization. Softmax classifiers exhibit visually distinct clusters in t-SNE but lose global structure in PCA/LDA, reflecting the absence of explicit latent variance modeling. Overall, the KL-regularized ResNet backbone consistently preserves richer geometric structure, which aligns with its stronger OOD performance.
(a) ResNet ZClassifier (t-SNE)
(b) VGG ZClassifier (t-SNE)
(c) NoKL ZClassifier (t-SNE)
(d) Softmax Classifier (t-SNE)
(e) ResNet ZClassifier (PCA)
(f) VGG ZClassifier (PCA)
(g) NoKL ZClassifier (PCA)
(h) Softmax Classifier (PCA)
(i) ResNet ZClassifier (LDA)
(j) VGG ZClassifier (LDA)
(k) NoKL ZClassifier (LDA)
(l) Softmax Classifier (LDA)
(m) ResNet ZClassifier (GMM)
(n) VGG ZClassifier (GMM)
(o) NoKL ZClassifier (GMM)
(p) Softmax Classifier (GMM)
Figure 4: CIFAR-100 latent logit visualizations across four classifiers and four projection methods: t-SNE, PCA, LDA, and GMM ellipses over PCA space. Colors denote class labels (10-class subset for clarity in GMM plots).
Calibration Robustness.
Figure 5 and Table 5 summarize the robustness of each classifier under additive Gaussian logit noise on CIFAR-100. Compared to CIFAR-10, accuracy declines more steeply across all models, reflecting the greater difficulty of the 100-class setting. ResNet ZClassifier remains the most stable, losing only $\sim$15% accuracy at $\text{STD}=1.0$, whereas the same configuration on CIFAR-10 degraded by merely $\sim$5%. VGG ZClassifier degrades more gradually than Softmax and NoKL, suggesting that KL-regularization aids calibration robustness even with a weaker backbone. Softmax and NoKL variants exhibit rapid degradation, with NoKL falling below 0.5 accuracy by $\text{STD}=1.0$.
Figure 5: CIFAR-100 calibration accuracy under additive Gaussian logit noise.
Table 5: CIFAR-100 calibration robustness (accuracy) under increasing logit noise standard deviation.
Model 0.00 0.22 0.44 0.67 0.89 1.11 1.33 1.56 1.78 2.00
ResNet ZClassifier 0.7370 0.7310 0.7280 0.7240 0.7180 0.7110 0.7050 0.6980 0.6900 0.6760
VGG ZClassifier 0.6220 0.6210 0.6190 0.5980 0.5780 0.5630 0.5480 0.5310 0.5150 0.4780
Softmax Classifier 0.7320 0.7200 0.6780 0.5970 0.4980 0.4250 0.3670 0.2890 0.3200 0.2680
NoKL ZClassifier 0.7000 0.6820 0.6350 0.5010 0.4230 0.3550 0.2940 0.2550 0.2350 0.2080
OOD Detection.
Table 6 presents the CIFAR-100 OOD detection results using KL divergence scores between the predicted Gaussian logits and the standard normal prior. Both ResNet- and VGG-based ZClassifiers achieve nearly perfect separability for synthetic OOD datasets (Gaussian and Uniform noise), with AUROC values close to 1.0 and FPR@95 near zero, indicating that latent Gaussian regularization is highly effective for large distributional shifts. For natural OOD datasets such as SVHN and CIFAR-10, performance moderately degrades, reflecting the greater difficulty of detecting semantically related domain shifts. The ResNet backbone consistently outperforms VGG, especially on SVHN (AUROC = 0.9682 vs. 0.9997), suggesting that deeper residual connections preserve discriminative latent structure under natural shifts. In stark contrast, the NoKL variant collapses across all OOD types, yielding AUROC $\approx 0.0$ and FPR@95 of 1.0, which confirms that KL regularization is critical for maintaining OOD separability in the latent space.
Figure 6 visualizes the KL score distributions for each (model, OOD dataset) combination. For Gaussian and Uniform noise, the in-distribution (blue) and OOD (orange) distributions are completely non-overlapping, consistent with the near-perfect detection metrics in Table 6. Natural shifts (CIFAR-10, SVHN) exhibit greater overlap between the distributions, particularly in the VGG variant, which explains the reduced AUROC in these settings. NoKL variants show severe distributional collapse, with almost complete overlap of ID and OOD score histograms, further illustrating the necessity of KL regularization for effective OOD detection.
(a) CIFAR-10 (ResNet)
(b) SVHN (ResNet)
(c) Gaussian (ResNet)
(d) Uniform (ResNet)
(e) CIFAR-10 (VGG)
(f) SVHN (VGG)
(g) Gaussian (VGG)
(h) Uniform (VGG)
(i) CIFAR-10 (NoKL)
(j) SVHN (NoKL)
(k) Gaussian (NoKL)
(l) Uniform (NoKL)
Figure 6: KL divergence score distributions for CIFAR-100 in-distribution vs. OOD datasets using ResNet, VGG, and NoKL ZClassifier variants.
Table 6: CIFAR-100 OOD detection via KL divergence.
Model OOD Dataset AUROC AUPR FPR@95
ResNet ZClassifier CIFAR-10 1.0000 1.0000 0.0000
ResNet ZClassifier SVHN 0.9682 0.8921 0.1063
ResNet ZClassifier Gaussian 1.0000 1.0000 0.0000
ResNet ZClassifier Uniform 1.0000 1.0000 0.0000
VGG ZClassifier CIFAR-10 1.0000 1.0000 0.0000
VGG ZClassifier SVHN 0.9997 0.9997 0.0000
VGG ZClassifier Gaussian 1.0000 1.0000 0.0000
VGG ZClassifier Uniform 0.9897 0.9835 0.0331
NoKL CIFAR-10 0.0000 0.1538 1.0000
NoKL SVHN 0.0000 0.3069 1.0000
NoKL Gaussian 0.0000 0.3069 1.0000
NoKL Uniform 0.0000 0.3069 1.0000